# FactTrack: Time-Aware World State Tracking in Story Outlines

### v1

@@ Appendix D Using FactTrack to Enhance Outline Generation
Table 16: The prompt given to GPT-4-Turbo to inject facts to a given event
@@
Table 17: The prompt given to GPT-4-Turbo to retrieve events to make a given event more consistent
Appendix E More Examples
We show three more correct examples of contradictions we detected in a story outline, in Table 18, Table 19, Table 20 and Table 21.
@@ Appendix E More Examples
Table 18: Extra Example 1. The contradiction occurs because in Event 1.2.1, unusual side effects have already been discovered and attributed to the building’s energy field, but in Event 2, it is implied that the causal relationship between the building and the side effects has not yet been discovered.
@@
Table 19: Extra Example 2. Although these two events were generated under one query of LLM, and there is still a factual contradiction. Event 2.3.1 indicates that the protagonist has completed the final challenge, but Event 2.3.2 suggests his mentor is helping him with the final challenge, creating a contradiction."
@@
Table 20: Extra Example 3. The contradiction arises because while Echo has already uncovered a conspiracy involving Dr. Kim and the government after event 2.3, but in event 3.2.2 there’s an implication that Echo is unaware of the government’s involvement in Dr. Kim’s work.
@@
Table 21: Extra Example 4. The contradiction arises in event 2.3, where Marcus and Leon have already reconciled, compared to event 3, where they just “begin to question the reasons behind their conflict,” indicating they are still in conflict.
@@
Appendix F Error Analysis
@@
Table 22, 23 and 22 illustrate three examples of such failures.

### v2

@@ Appendix E Using FactTrack to Enhance Outline Generation
Table 16: The prompt given to GPT-4 to inject facts to a given event
@@
Table 17: The prompt given to GPT-4 to retrieve events to make a given event more consistent
Appendix F More Examples
We show five correct examples of contradictions we detected in a story outline in Table 18, Table 19, Table 20, Table 21 and 22. We can see that our approach performs detailed comparisons at a fine-grained fact level rather than simply checking for contextual continuity.
@@ Appendix F More Examples
Table 18: Example 1. The contradiction occurs because in Event 1.2.1, unusual side effects have already been discovered and attributed to the building’s energy field, but in Event 2, it is implied that the causal relationship between the building and the side effects has not yet been discovered.
@@
Table 19: Example 2. Although these two events were generated under one query of LLM, and there is still a factual contradiction. Event 2.3.1 indicates that the protagonist has completed the final challenge, but Event 2.3.2 suggests his mentor is helping him with the final challenge, creating a contradiction."
@@
Table 20: Example 3. The contradiction arises because while Echo has already uncovered a conspiracy involving Dr. Kim and the government after event 2.3, but in event 3.2.2 there’s an implication that Echo is unaware of the government’s involvement in Dr. Kim’s work.
@@
Table 21: Example 4. The contradiction arises in event 2.3, where Marcus and Leon have already reconciled, compared to event 3, where they just “begin to question the reasons behind their conflict,” indicating they are still in conflict.
Event Pair
Atomic Fact Pairs Detected as Contradict
Table 22: Example 5. The contradiction arises because in event 1.3.3, the group already gains knowledge of hidden secrets that are the reason they have been brought together, which contradicts the indication that before event 3.2, they are unaware about those secrets.
@@
Appendix G Error Analysis
@@
Table 23, 24 and 23 illustrate three examples of such failures.
Appendix H Usage of AI Assistants
During the development process, we drafted the code framework and data structures by ourselves initially, using Copilot to assist in enhancing development efficiency. Additionally, in the writing phase, we only employed ChatGPT for proofreading purposes, such as correcting grammatical errors and properly formatting tables.
Appendix I Model Size and Computation Budget
We utilized the LLaMA2-7B-Chat model, a 7 billion parameter generative language model. Additionally, the NLI and retrieval models used were both under 1 billion parameters. The total computational cost for all experiments did not exceed 100 GPU hours on an A6000, and the cost of API of GPT-series was kept under \$500 USD.